Use Polya’s four-step problem-solving strategy and the problem-

solving procedures to solve each of the following exercises.


(a) An android phone and a Bluetooth speaker together cost Php6500.00

. The phone costs Php2000.00 more than the speaker. What is

the cost of the phone?

(b) In how many ways can you answer a 12-question true-false test

if you answer each question with either a “true” or a “false”?

(c) A total of 364 first-year students of CMU are currently enrolled

in the subject GEC14. If there are 26 more girls than boys, how

many girls are there?


1
Expert's answer
2021-10-21T13:15:06-0400

a)

x- cost of android phone,

y - cost of Bluetooth speaker


"x+y=6500"

"x=2000+y"


"2000+2y=6500"

"y=2250"

"x=2000+2250=4450"


Answer: the phone costs Php 4450.


b)

"N=2^{12}=4096" ways


c)

x - number of girls,

y - number of boys


"x+y=364"

"x=y+26"


"2y+26=364"

"y=338\/2=169"

"x=169+26=195"


Answer: 195 girls